Shift report 20000903 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Sep 03 2000 - 08:00:31 EDT

  • Next message: Yury Blyakhman: "Re: BB ADC gain and multiplicity calculation"

    ------------------------------------------------------------
    Date: Sun Sep  3 08:00:31 2000, Shift: 00:00-08:00
    Supervisor: Hiro, with: Michael
    Log book#5 pages 83 - 85
    
    Shift Summary:
    ------------------------------------------------------------
    1) Beam from 3am until now intensity slowly dropped.
    Blue was 1.5E10 no 1.0E10 Yellow was 1.9E10 now 1.5E10
    
    2) Took runs 2516, 2517 and 2518 with MRS at 40 degrees
    FF at 4 deg BF at 4 deg all magnets at B setting
    Run 2518 is continueing. We took about 700 000 trigger 1
    events
    
    3) BB Right Module 9 looks dead, (this was also  true 
    yesterday afternoon.)
    
    One argon bottle is empty. T3 plane 3 U1 has a gap from
    channels 16-32. While T3 plane 3 U2 has a lot of extra
    hits for channels 16-32.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Sep 03 2000 - 08:00:50 EDT